i am thinking of a 2 digit number. the digit in the tens place is greater than the digit in the ones place. the sum of its digits is 12. the product of its digits is 35. what number is thinking of?
Answer by math_helper(2461)

Let x be the 10's digit, and y the one's digit.
x + y = 12
xy = 35
x > y
x = 35/y —> x + y = 35/y + y = 12
35 + y^2 = 12y
y^2 - 12y + 35 = 0
(y-7)(y-5) = 0
y=7 or y=5
y=7 —> x=5, no good because we only want solutions where x>y
y=5 —> x=7 (ok, xy=35, x+y=12, and x>y)

Ans: The number is 75.